A furniture shop refinishes chairs. Employees use two methods to refinish a chair. method I takes 1 hour and costs $3. Method II takes 1 1/2 hours and material costs $1.50. Last Week, they took 36 hours and spent $60 refinishing chairs. How many did they refinish with each method?
Answer by stanbon(75887)

A furniture shop refinishes chairs. Employees use two methods to refinish a chair. method I takes 1 hour and costs $3. Method II takes 1 1/2 hours and material costs $1.50. Last Week, they took 36 hours and spent $60 refinishing chairs. How many did they refinish with each method?
--------------
Let Method I be A; Let Method II be B.
1st: A+(3/2)B=36
2nd: 3A+(3/2)B=60
Subtract 1st from 2nd to get:
2A=24
A=12 hours (Time they used Method I)
B=24 hours (Time they used Method II)
But how many chairs did they finish with each method?
In 12 hours Method A will finish 12 chairs.
In 24 hours Method B will finish 24/(3/2)= 16 chairs.
Cheers,
Stan H.